OpenStudy (teller):

Can someone make sure I did this right? Suppose that the functions r and s are defined for all real numbers x as follows. r(x)=x+6 s(x)=2x-4 Write the expressions for (s+r)(x) and (s-r)(x) and evaluate (s.r)(1) (s+r)(x)=(2x-4)+(x+6) s-r)(x)=(2x-4)-(x+6) (s.r)(1)= 2(x-2) over x+6

OpenStudy (anonymous):

right but you are not done

OpenStudy (anonymous):

for the first two, combine like terms

OpenStudy (teller):

So it would be (s+r)(x)= (2x-x)+(4+6) (s-r)(x)=2x-x)-(4+6)

OpenStudy (teller):

For the last part I know I got wrong because my thing doesn't have a fraction sign so I know that that can't be my answer.

OpenStudy (anonymous):

\[(s+r)(x)=(2x-4)+(x+6)=3x+2\]

OpenStudy (teller):

what's that for ?

OpenStudy (teller):

where did you get the 2x+2?

OpenStudy (anonymous):

that is what you get when you combine like terms

OpenStudy (teller):

but the problem has a divide symbol .. not a plus

OpenStudy (teller):

wouldn't you divide?

OpenStudy (anonymous):

i copied and pasted what you wrote

OpenStudy (anonymous):

for number 1

OpenStudy (teller):

I'm confused... so that would be for the first one?

OpenStudy (anonymous):

yes

OpenStudy (teller):

So the second one would be (s+r)(x)=(2x−4)-(x+6)=x-10?

OpenStudy (teller):

so what would the last choice be ?